cosecA + cosec2A + cosec4A + cosec 8A …. n terms =

Grade:10

5 Answers

Arun Kumar IIT Delhi
9 years ago
If its


Now its doesn't seem to solve into anything.


Arun Kumar
Sanip Panta
7 years ago
The answer of this question is: CotA-cotnA °n is the angle of last term as cosecnA. For example:CosecA+cosec2A+cosec4A+cosec8A=cotA-cot8A
Sanip Panta
7 years ago
The answer was a bit mistake. Actually it was cotA/2 - cotnA #bA is the angle of last term. For example :CosecA+cosec2A +cosec4A +cosec8A +cosec16A = cotA/2 - cot16A
Shree Ram Gupta
6 years ago
= cotA - cot8A= cotA - cos8A/sin8A= cotA - 2cos^2(4A) - 1/2sin4Acos4A= cotA - 2cos^2(4A)/2sin4Acos4A + 1/2sin4Acos4A= cotA - cos4A/sin4A + 1/sin8A= cotA - 2cos^2(2A) - 1/2sin2Acos2A + cosec8A= cotA - 2cos^2(A)/2sin2Acos2A + 1/2sin2Acos2A + cosec8A= cotA - cos2A/sin2A+ 1/sin4A+cosec8AcotA - 2cos^2(A) - 1/2sinAcosA + cosec4A + cosec8A=cosA/sinA - cosA/sinA + 1/sin2A + cosec4A + cosec8A= cosec2A + cosec4A + cosec8A
Ajay Saini
2 years ago
We know that cotA-cot2A=cosec2A then
CosecA + cosec2A +cosec3A.......cosec2^n-1
= cota/2-cota+cotA-cot2A+cot2A-cot3A.....cot2^n-2+cot2^n-1
=Cota/2-cot2^n-1